OpenStudy (anonymous):

How do we prove that subtracting an even number by 2 still gives an even number?

OpenStudy (anonymous):

I asked a similar question earlier and based on that have made an attempt similar tot he answer I was given. Can somebody check that I am right?...

OpenStudy (anonymous):

Lets say that the starting even number is x. x = 2k where k can be any integer, even or odd. x - 2 = 2k - 2 x - 2 = 2(k-1) As this is still dividable by 2, it is an even number still

OpenStudy (anonymous):

Am I correct?

OpenStudy (mary.rojas):

10-2=8 8 is even is that what you mean?

OpenStudy (anonymous):

well yes, but algebraic proof for all even numbers

OpenStudy (anonymous):

Proving that [Even number 1] - 2 = [even number 2] i.e. it will never be odd

jhonyy9 (jhonyy9):

so you know that every even has a form of 2n and 2 is even too so 2n-2 = 2(n-1) so note (n-1=k than 2(n-1)=2k what mean that 2k will be allways even hope that you understand it now sure

OpenStudy (mary.rojas):

oh sorry lol, am not sure but good luck

jhonyy9 (jhonyy9):

and yes you are right

OpenStudy (anonymous):

So my proof was right?

OpenStudy (anonymous):

Thanks :)

jhonyy9 (jhonyy9):

good luck

jhonyy9 (jhonyy9):

my pleasure